# Observation of thermal microwave photons with a Josephson junction detector

### v1

@@ Abstract
In this work we describe and operate a microwave single photon detector used to observe the ticking and the statistics of thermal photons emitted by a resonator at ultracryogenic temperatures. The Josephson junction-based sensor shows sub-Hz dark count rate and high efficiency, combined with a reduced operation complexity compared to qubit-based designs [43, 9]. The SPD was used to observe single microwave photons emitted by a single-mode and multi-mode thermal source, i. e. a copper cavity. Although widely accepted [18, 59, 60], an experimental demonstration of the super-Poissonian statistics of thermal microwave photons was, to our knowledge, missing. The measured photon rate is consistent with thermal emission from the microwave cavity modes, and closely follows the expected temperature dependence. The photon distribution displays a clear reduction of the mean to standard deviation ratio when the light is emitted mainly by one mode, which instead remains unitary for dark count and multi-mode emission. The dynamical range of the detector extends from its dark count rate to the kHz range and is currently limited by the acquisition electronics’ speed. The detector dynamics was not further investigated, and will be the subject of future works. The SPD efficiency can be better characterised with an on-demand microwave photon source [61, 62], to disentangle it from other systematic uncertainties. Future efforts to improve the SPD aim mainly at improving its dark count rate. For instance, a junction with lower $I_{C}$ could enter the phase diffusion regime, which is expected to increase its lifetime, and thereby reduce the dark count rate [17]. Advancing the screening and filtering of the setup is also foreseen to improve the detector performances. In perspective, this work paves the way to a deeper examination of thermal light’s quantum properties, and to their use to study quantum chaos [21], ghost imaging [63, 64, 65], and more [18]. The device itself finds numerous applications in the field of precision physics beyond the Standard Model, as for instance Axion searches [22, 51]. On one hand, the performance requirements needed to drastically improve current Axion experiments are already met by the current device, on the other hand, these apparata typically include strong magnetic field and long operation times, posing a challenge to the screening and long-term-stability of the detector. Nevertheless, the implementation of single microwave photon detectors such as the one presented here fosters the emergence of next-generation beyond the Standard Model experiments.
@@ Acknowledgments
The work is supported by the Russian Science Foundation (Project No. 19-79-10170). This project is also supported by the Italian Institute of Nuclear Physics (INFN) within the QUAX experiment. The authors wish to thank G. Carugno for triggering the interest on photon detectors and for the multiple advice in the course of this work, A.A. Yablokov for writing automated Python-based software, and I.V. Rakut’ for sample holder fabrication.
@@
The general idea of the work was suggested by L.S.K., N.C., A.V.G. and A.L.P. The sample fabrication was performed by A.V.G and A.V.B. A.L.P. ran the experiment with support of A.V.G., A.V.C. and L.S.R. Data were analyzed by N.C., A.L.P., A.V.G. and A.V.C. A.L.P., A.V.G., L.S.R. and A.V.B. designed and assembled the experiment. The cavity was designed by N.C., and tested by N.C., A.L.P., A.V.G. and A.V.C. All authors, led by N.C. and A.L.P., contributed to the manuscript with figures created by N.C., A.V.G. and A.V.C.

### v2

@@ Discussion
In this work we describe and operate a microwave single photon detector used to observe the clicking and the statistics of thermal photons emitted by a resonator at ultracryogenic temperatures. The Josephson junction-based sensor shows sub-Hz dark count rate and high efficiency, combined with a reduced operation complexity compared to qubit-based designs [42, 9]. The SPD was used to observe single microwave photons emitted by a single-mode and multi-mode thermal source, i. e. a copper cavity. Although widely accepted [18, 59, 60], an experimental demonstration of the super-Poissonian statistics of thermal microwave photons was, to our knowledge, missing. The measured photon rate is consistent with thermal emission from the microwave cavity modes, and closely follows the expected temperature dependence. The photon distribution displays a clear reduction of the mean to standard deviation ratio when the light is emitted mainly by one mode, which instead remains unitary for dark count and multi-mode emission. The dynamical range of the detector extends from its dark count rate to the kHz range and is currently limited by the acquisition electronics’ speed. The detector dynamics was not further investigated, and will be the subject of future works. The SPD efficiency can be better characterised with an on-demand microwave photon source [61, 62, 63, 64], to disentangle it from other systematic uncertainties. Future efforts to improve the SPD aim mainly at improving its dark count rate. For instance, a junction with lower $I_{C}$ could enter the phase diffusion regime, which is expected to increase its lifetime, and thereby reduce the dark count rate [17]. Advancing the screening and filtering of the setup is also foreseen to improve the detector performances. In perspective, this work paves the way to a deeper examination of thermal light’s quantum properties, and to their use to study quantum chaos [20], ghost imaging [65, 66, 67], and more [18]. The device itself finds numerous applications in the field of precision physics beyond the Standard Model, as for instance Axion searches [21, 50]. On the one hand, the performance requirements needed to drastically improve current Axion experiments are already met by the current device, on the other hand, these apparata typically include strong magnetic field and long operation times, posing a challenge to the screening and long-term-stability of the detector. Nevertheless, the implementation of single microwave photon detectors such as the one presented here fosters the emergence of next-generation beyond the Standard Model experiments.
@@ Acknowledgments
The work is supported by the Russian Science Foundation (Project No. 19-79-10170). This project is also supported by the Italian Institute of Nuclear Physics (INFN) within the QUAX experiment. The authors wish to thank G. Carugno for triggering the interest on photon detectors and for the multiple advice in the course of this work, A.A. Yablokov for writing an automated Python-based software, and I.V. Rakut’ for sample holder fabrication.
@@
The general idea of the work was suggested by L.S.K., N.C., A.V.G. and A.L.P. The sample fabrication was performed by A.V.G and A.V.B. A.L.P. ran the experiment with support of A.V.G., A.V.C. and L.S.R. Data were analyzed by N.C., A.L.P., A.V.G. and A.V.C. A.L.P., A.V.G., L.S.R. and A.V.B. designed and assembled the experiment. The cavity was designed by N.C., and tested by N.C., A.L.P., A.V.G. and A.V.C. All authors, led by N.C. and A.L.P., contributed to manuscript with figures created by N.C., A.V.G. and A.V.C.